The sum of the positive odd integers less than 50 is subtracted from the sum of the positive even integers less than or equal to 50. What is the resulting difference?
Answer by Fombitz(32388)

(2+4+6+...+48+50)-(1+3+5+...+47+49)=(2-1)+(4-3)+(6-5)+...+(48-47)+(50-48)
(2+4+6+...+48+50)-(1+3+5+...+47+49)=(1)+(1)+(1)+...+(1)+(1)<--- 25 pairs
(2+4+6+...+48+50)-(1+3+5+...+47+49)=25